# CHEMOUT: CHEMical complexity in star-forming regions of the OUTer Galaxy

### v1

@@ Abstract
Aims. To understand how chemistry evolves across the Milky Way, deriving molecular abundances in star-forming regions in the outer Galaxy is essential to refine chemical models designed for environments with sub-Solar metallicity.
Methods. We analyzed IRAM 30m observations in several spectral windows at 3 and 2 mm, towards a sample of 35 sources located at Galactocentric distances $\sim 9-24$ kpc in the context of the project ’CHEMical complexity in star-forming regions of the outer Galaxy’ (CHEMOUT).
@@
Conclusions. These results suggest that, in the outer Galaxy, the efficiency in the formation of most of the molecules studied, scaling with the availability of the parent element, is at least as high as in the local Galaxy, or even higher (e.g., for H13CO+). Yet, other species, as SO, are characterized by a lower formation efficiency. These outcomes have important implications for the chemistry occurring in the outermost star-forming regions of the Galaxy and can help to constrain models adapted for lower metallicity environments.
@@ 1 Introduction
The outer Galaxy (OG), namely the portion of the Galactic disk beyond the Solar circle, is located at Galactocentric distances ($R_{\text{GC}}$) approximately between 9 and 24 kpc, and shows chemical properties significantly different from those of the inner Galaxy. Several chemical Galactocentric gradient studies have been conducted, mapping the differences in elemental composition between the inner and outer Milky Way (e.g., C, N, O, Fe), using various tracers (i.e., H II regions, classical Cepheids, young OB stars, and planetary nebulae, among others). These works, focused on observational trends (e.g., Luck 2018; Bragança et al. 2019; Méndez-Delgado et al. 2022) and chemodynamical studies (Jacobson et al. 2016; Stanghellini & Haywood 2018; Palla et al. 2020; Santos-Peral et al. 2021), indicate that the overall metallicity in the OG is lower than the Solar one (e.g., a factor of four lower at $R_{\text{GC}}=19\penalty 10000\$kpc; Shimonishi et al. 2021). The elemental abundances of carbon, oxygen, and nitrogen - the three most abundant elements in the universe after hydrogen and helium - decrease as a function of $R_{\text{GC}}$ (e.g., Esteban et al. 2017; Arellano-Córdova et al. 2020), similarly to other heavier elements, following the radial metallicity gradient. This observational result is linked to the traditional view of the OG as being poorer in molecules than the Solar neighborhoods and the inner Galaxy. Moreover, based on the aforementioned scarcity of elements crucial for the formation of complex molecules and (rocky) planets, chemical evolution models excluded most of the OG from the so-called Galactic habitable zone (GHZ), that is, the region in the Milky Way where the probability for the development of habitable planets is highest: the optimal region is predicted to be an annulus of the Galactic disk from $\sim 7\penalty 10000\$kpc extended up to $R_{\text{GC}}\sim 9\penalty 10000\$kpc, with maximum habitability at $\sim 8$ kpc (Spitoni et al. 2014, 2017). This led to the consideration of the OG not being suitable for the formation of planetary systems in which Earth-like planets could be born and might be capable of sustaining life (Prantzos 2008; Ramírez et al. 2010).
@@
Recent observational results challenged this prediction. First, the occurrence of Earth-like planets does not seem to depend on metallicity (e.g., Maliuk & Budaj 2020; Mulders et al. 2018). Second, the low rate of disruptive events in the OG suggests that it might be more favorable than the inner Galaxy for preserving life on potentially habitable planets (e.g., Piran & Jimenez 2014; Vukotić et al. 2016). Third, the detection of several complex organic molecules (COMs; e.g., CH3OH, CH3CCH, CH3OCHO, CH3OCH3) of prebiotic interest in low metallicity environments, such as star-forming regions in the OG (Shimonishi et al. 2021; Bernal et al. 2021; Fontani et al. 2022, 2024; Koelemay & Ziurys 2025), and the Magellanic Clouds (Sewiło et al. 2018, 2022; Shimonishi et al. 2023; Golshan et al. 2024). Overall, these results suggest that environments prone to the presence of molecular precursors of life and Earth-like planets might be widespread in the Milky Way, including also the OG.
@@
In this line, the project ’CHEMical complexity in the outer Galaxy’ (CHEMOUT; Fontani et al. 2022, 2022b, Colzi et al. 2022, Fontani et al. 2024), performed with the IRAM 30m single-dish telescope, is aimed at deriving the molecular composition in 35 star-forming regions of the OG ($R_{\text{GC}}\sim 9-24\penalty 10000\$kpc). The main results obtained in the previous works of the CHEMOUT project are: (1) both the molecular detection rates and the molecular abundances of some organic species, namely HCO, H2CO, and CH3OH, do not vary significantly with $R_{\text{GC}}$ (Fontani et al. 2022, 2022b); (2) the 14N/15N ratio as a function of $R_{\text{GC}}$ is consistent with Chemical Evolution models in which 15N comes from novae (Colzi et al. 2022); (3) the high-angular resolution maps of the target with the highest $R_{\text{GC}}$ (WB89-670 at $R_{\text{GC}}\sim 23.4$ kpc) reveal chemical differentiation at core scales and, above all, a [C/H] abundance higher than expected at such high Galactocentric distance (Fontani et al. 2024).
@@
The results obtained so far indicate a lack of decrease in the inspected molecular abundances, which, however, was limited to three species (HCO, H2CO, and CH3OH) and only 15 sources. In this paper, we analyze the molecules detected in more than $60\%$ of the CHEMOUT sample according to Fontani et al. (2022), namely HCN, HCO+, c-C3H2, H13CO+, HCO, and SO, and revise the molecular abundances of H2CO and CH3OH from Fontani et al. (2022b), using H2 column densities from the new NIKA2 dust continuum maps. In Sect. 2, we describe the observational data used in this paper. In Sect. 3, we describe the methods implemented in the data analysis. The results are presented and discussed in Sect. 4 and 5. The conclusions and the outlooks are summarized in Sect. 6.
@@ 2.1 IRAM 30m observations
This work is based on the observations done for the CHEMOUT project, described in Papers I and II (Fontani et al. 2022, 2022b), performed with the Institut de RadioAstronomie Millimétrique (IRAM) 30m telescope. The sample of sources analyzed in this paper is composed of the 35 star-forming regions, presented in Table 1 of Fontani et al. (2022). In some spectra, two velocity features are identified, and the sources listed in Table 5 in this paper include both velocity components. In the analysis conducted, we used spectral windows at 3 and 2 mm, collected with the EMIR receiver, and described in Table 1. The spectral windows of the project 004-18 (see Table 1) have been optimized to observe, at 3 mm, the J=10J=1-0 transitions of 13C and 15N isotopologues of HCN and HNC and, at 2 mm, the J=21J=2-1 line of DNC. The spectral windows observed in the project 042-21, at 3 and 2 mm, have been designed to observe several methanol and formaldehyde lines. For further details (e.g., pointing and focus, and weather conditions), we refer to Fontani et al. (2022, 2022b).
@@ 2.2 NIKA2 bolometer observations
Millimeter continuum emission of 31 out of the 35 CHEMOUT targets was observed with the New IRAM KID Arrays 2 (NIKA2; Perotto et al. 2020) bolometer of the IRAM 30m telescope in Pico Veleta, Spain. Observations were taken during the winter 2021 and 2023 pool seasons (February). The observations cover an almost circular region of radius $\sim 9$ arcmin, centered at the positions where the IRAM 30m spectra have been observed (Table 5). The dual-band capability of NIKA2 allowed us to simultaneously observe the dust thermal continuum emission at 2.0 and 1.2 mm (150 GHz and 260 GHz, respectively). The sensitivity achieved was $\sim 1.5$ mJy/beam at 1.2 mm and 0.5 mJy/beam at 2.0 mm. A more complete description of the observations and a detailed analysis of the maps will be presented in a forthcoming paper of the CHEMOUT series (Fontani et al. in prep). Here, we have used the maps only to derive the H2 column densities (Sect. 3.2) to compute the molecular fractional abundances with respect to H2 (Sect. 4.2).
Table 1: Spectral windows at 3 mm and 2 mm wavelength observed in the CHEMOUT project, used in this paper, and the related observational parameters. [^1] The observational parameters presented includes velocity resolution ($V_{\text{res}}$) of the spectra, main beam efficiency ($B{{}_{\text{eff}}}$), forward efficiency ($F_{\text{eff}}$), and system temperature ($T_{\text{sys}}$).
@@ 3.1 Line fitting
The spectra were fitted using the MAdrid Data CUBe Analysis (MADCUBA [^2] MADCUBA is a software developed in the Madrid Center of Astrobiology (CAB, CSIC-INTA) which enables the visualization and analysis of single spectra and data cubes: https://cab.inta-csic.es/madcuba/; Martín et al. 2019) software. The transitions were identified using the Spectral Line Identification and LTE Modelling (SLIM) tool of MADCUBA, which makes use of the Jet Propulsion Laboratory (JPL; Pickett et al. 1998) and Cologne Database for Molecular Spectroscopy (CDMS; Müller et al. 2001) catalogs. The lines were fitted using the AUTOFIT function of SLIM. This function produces the Gaussian synthetic spectrum that best matches the observed spectrum, assuming local thermodynamic equilibrium (LTE) conditions and the same fitting input parameters for all transitions of a given molecular species, which are: total column density $N_{\text{tot}}$, excitation temperature $T_{\text{ex}}$, radial systemic velocity of the source $V_{\text{LSR}}$, full-width at half-maximum of the line FWHM, and angular size of the emission $\theta_{\text{S}}$.
@@
For the source size, we assumed that the molecular emission fills the telescope beam, except for methanol and formaldehyde, where an estimation of the emission size has been made in Fontani et al. (2022b) and updated in this paper as discussed in Sect. 4.1.1. Such an assumption is justified by the spatial extent of the emission of the dust (see Figs. 4, 5, 6, 7), as well as by the updated H2CO and CH3OH emission sizes shown in Table 8. The excitation temperature, whenever an estimate was not possible through the fit, has been fixed to the one shown in Table 5, estimated from the methanol lines for the 15 targets observed in Fontani et al. (2022b). For the remaining sources, we adopted an average $T_{\text{ex}}$ of 10.7 K, derived from those sources for which methanol transitions were available. This approximation was necessary because it was not possible to estimate the temperature individually for each source based on the observed transitions. To evaluate the impact of this assumption, we computed $N_{\text{tot}}$ in the range of excitation temperatures derived from methanol ($7\leq T_{\text{ex}}\leq 15\penalty 10000\ K$; see Table 5). In this $T_{\text{ex}}$ range, the densities vary at most by a factor of 2 from the values shown in Tables 7 and 8, for the species studied (lowest for HCN, HCO+, and H13CO+, and highest for SO).
@@
As a detection criterion, we have considered a line as detected if the signal-to-noise ratio (SNR) is greater than 3 (an estimation of the rms noise is given in Fontani et al. 2022), while the detection is considered as tentative for $2.5\leq\text{SNR}\leq 3$. If a line is undetected (e.g., SNR $\leq 2.5$) MADCUBA provides a $3\sigma$ upper limit estimate to the line integrated intensity from Eq. (29) of Martín et al. (2019). The upper limit on the integrated intensity is converted into a column density upper limit using the SLIM parameters $T_{\text{ex}}$ and $\theta_{\text{S}}$. For these estimates, the average FWHM for each species across the sample was adopted (Sect. 5.2). This approach assumes that different molecules trace gases with different physical, kinematic, and excitation conditions.
@@
The total column densities derived from transitions observed at frequencies corresponding to beam sizes different from 28′′ (i.e., the largest beam size among the IRAM 30m telescope spectral windows) have been rescaled by a factor of $(\frac{\theta_{\text{beam}}}{28^{\prime\prime}})^{2}$. For H2CO and CH3OH, the column densities have been estimated assuming a specific angular emission size (see Table 8, and Sect. 4.1.1). In cases where the emission size exceeds 28′′, no correction was applied, under the assumption of uniform emission within the beam. Conversely, when the source size is smaller than 28′′, the same scaling factor was applied to account for beam dilution effects. The estimation of the molecular fractional abundances has been performed using:
@@ 3.2 Continuum analysis
For this paper, the analysis of these maps is limited to the extraction of the flux density from an angular region of $28^{\prime\prime}$ centered on the pointing centers of the molecular spectroscopic data. The dimension of the extracted region has been decided to match the larger beam size of IRAM 30m in the studied spectral windows (Table 1). A complete description and a deeper analysis of the NIKA2 observations will be given in a forthcoming paper (Fontani et al. in prep.).
The column densities of H2, $N_{\text{H}_{2}}$ have been derived from the dust continuum emission maps. In the scales (and densities) traced by our observations, the thermal dust emission is expected to be optically thin. Under this assumption, we can use the following equation to derive $N_{\text{H}_{2}}$ from the continuum flux density (e.g., Battersby et al. 2011):

### v2

@@ Abstract
Aims. To understand how chemistry evolves across the Milky Way, deriving molecular abundances in star-forming regions in the outer Galaxy is essential to refining chemical models designed for environments with subsolar metallicity.
Methods. We analyzed IRAM 30 m observations in several spectral windows at 3 and 2 mm, toward a sample of 35 sources located at galactocentric distances of $\sim 9-24$ kpc in the context of the project CHEMical complexity in star-forming regions of the outer Galaxy (CHEMOUT).
@@
Conclusions. These results suggest that, in the outer Galaxy, the efficiency in the formation of most of the molecules studied, scaling with the availability of the parent element, is at least as high as in the local Galaxy, or perhaps even higher (e.g., for H13CO+). Yet, other species, such as SO, are characterized by a lower formation efficiency. These outcomes have important implications for the chemistry occurring in the outermost star-forming regions of the Galaxy and can help to constrain models adapted to lower metallicity environments.
@@ 1 Introduction
The outer Galaxy (OG), namely the portion of the Galactic disk beyond the solar circle, is located at galactocentric distances ($R_{\text{GC}}$) approximately between 9 and 24 kpc, and it shows chemical properties significantly different from those of the inner Galaxy. Several chemical galactocentric gradient studies have been conducted, mapping the differences in elemental composition between the inner and outer Milky Way (e.g., C, N, O, Fe), using various tracers (i.e., H II regions, classical Cepheids, young OB stars, and planetary nebulae, among others). These works, focused on observational trends (e.g., luck2018cepheid; bragancca2019radial; mendez2022gradients) and chemodynamical studies (jacobson2016gaia; stanghellini2018galactic; palla2020chemical; santos2021ambre), indicate that the overall metallicity in the OG is lower than the solar one (e.g., a factor of four lower at $R_{\text{GC}}=19\penalty 10000\$kpc; shimonishi2021detection). The elemental abundances of carbon, oxygen, and nitrogen - the three most abundant elements in the Universe after hydrogen and helium - decrease as a function of $R_{\text{GC}}$ (e.g., esteban2017radial; arellano2020galactic), similarly to other heavier elements, following the radial metallicity gradient. This observational result is linked to the traditional view of the OG as being poorer in molecules than the solar neighborhoods and the inner Galaxy. Moreover, based on the aforementioned scarcity of elements crucial for the formation of complex molecules and (rocky) planets, chemical evolution models excluded most of the OG from the so-called Galactic habitable zone (GHZ); that is, the region in the Milky Way where the probability for the development of habitable planets is highest: the optimal region is predicted to be an annulus of the Galactic disk from $\sim 7\penalty 10000\$kpc extended up to $R_{\text{GC}}\sim 9\penalty 10000\$kpc, with maximum habitability at $\sim 8$ kpc (spitoni2014galactic; spitoni2017evolution). This led to the consideration of the OG not being suitable for the formation of planetary systems in which Earth-like planets could be born and might be capable of sustaining life (prantzos2008galactic; ramirez2010possible).
@@
Recent observational results challenged this prediction. First, the occurrence of Earth-like planets does not seem to depend on metallicity (e.g., maliuk2020spatial; mulders2018exoplanet). Second, the low rate of disruptive events in the OG suggests that it might be more favorable than the inner Galaxy for preserving life on potentially habitable planets (e.g., piran2014possible; vukotic2016grandeur). Third, the detection of several complex organic molecules (COMs; e.g., CH3OH, CH3CCH, CH3OCHO, CH3OCH3) of prebiotic interest in low-metallicity environments, such as star-forming regions in the OG (shimonishi2021detection; bernal2021methanol; fontani2022chemout; fontani2024chemout; koelemay2025no) and the Magellanic Clouds (sewilo2018detection; sewilo2022alma; shimonishi2023detection; golshan2024high). Overall, these results suggest that environments prone to the presence of molecular precursors of life and Earth-like planets might be widespread in the Milky Way, including the OG.
@@
The project CHEMical complexity in the outer Galaxy (CHEMOUT; fontani2022chemout fontani2022chemout, fontani2022chemoutII, colzi2022chemout colzi2022chemout, fontani2024chemout fontani2024chemout), performed with the IRAM 30 m single-dish telescope, is aimed at deriving the molecular composition in 35 star-forming regions of the OG ($R_{\text{GC}}\sim 9-24\penalty 10000\$kpc). The main results obtained in the previous works of the CHEMOUT project are (1) both the molecular detection rates and the molecular abundances of some organic species –namely HCO, H2CO, and CH3OH– do not vary significantly with $R_{\text{GC}}$ (fontani2022chemout fontani2022chemout, fontani2022chemoutII); (2) the 14N/15N ratio as a function of $R_{\text{GC}}$ is consistent with chemical-evolution models in which 15N comes from novae (colzi2022chemout); (3) the high-angular-resolution maps of the target with the highest $R_{\text{GC}}$ (WB89-670 at $R_{\text{GC}}\sim 23.4$ kpc) reveal chemical differentiation at core scales and, above all, a [C/H] abundance higher than expected at such a high galactocentric distance (fontani2024chemout).
@@
The results obtained so far indicate a lack of decrease in the inspected molecular abundances, which, however, was limited to three species (HCO, H2CO, and CH3OH) and only 15 sources. In this paper, we analyze the molecules detected in more than $60\%$ of the CHEMOUT sample according to fontani2022chemout, namely HCN, HCO+, c-C3H2, H13CO+, HCO, and SO, and revise the molecular abundances of H2CO and CH3OH from fontani2022chemoutII, using H2 column densities from the new NIKA2 dust continuum maps. In Sect. 2, we describe the observational data used in this paper. In Sect. 3, we describe the methods implemented in the data analysis. The results are presented and discussed in Sect. 4 and 5. The conclusions and the outlooks are summarized in Sect. 6.
@@ 2.1 IRAM 30 m observations
This work is based on the observations done for the CHEMOUT project, described in Papers I and II (fontani2022chemout fontani2022chemout, fontani2022chemoutII), performed with the Institut de RadioAstronomie Millimétrique (IRAM) 30 m telescope. The sample of sources analyzed in this paper is composed of the 35 star-forming regions, presented in Table 1 of fontani2022chemout. In some spectra, two velocity features are identified, and the sources listed in Table 5 in this paper include both velocity components. In the analysis conducted, we used spectral windows at 3 and 2 mm, collected with the EMIR receiver and described in Table 1. The spectral windows of the project 004-18 (see Table 1) have been optimized to observe, at 3 mm, the J=10J=1-0 transitions of 13C and 15N isotopologues of HCN and HNC and, at 2 mm, the J=21J=2-1 line of DNC. The spectral windows observed in the project 042-21, at 3 and 2 mm, have been designed to observe several methanol and formaldehyde lines. For further details (e.g., pointing and focus, and weather conditions), we refer the reader to fontani2022chemout (fontani2022chemout, fontani2022chemoutII).
@@ 2.2 NIKA2 bolometer observations
Millimeter continuum emission of 31 out of the 35 CHEMOUT targets was observed with the New IRAM KID Arrays 2 (NIKA2; perotto2020) bolometer of the IRAM 30 m telescope in Pico Veleta, Spain. Observations were taken during the winter 2021 and 2023 pool seasons (February). The observations cover an almost circular region with a radius of $\sim 9$ arcmin, centered at the positions where the IRAM 30 m spectra have been observed (Table 5). The dual-band capability of NIKA2 allowed us to simultaneously observe the dust’s thermal continuum emission at 2.0 and 1.2 mm (150 GHz and 260 GHz, respectively). The sensitivity achieved was $\sim 1.5$ mJy/beam at 1.2 mm and 0.5 mJy/beam at 2.0 mm. A more complete description of the observations and a detailed analysis of the maps will be presented in a forthcoming paper of the CHEMOUT series (Fontani et al. in prep). Here, we have used the maps only to derive the H2 column densities (Sect. 3.2) to compute the molecular fractional abundances with respect to H2 (Sect. 4.2).
Table 1: Observational set ups of CHEMOUT used in this paper. [^1] The observational parameters presented include the frequency ranges observed (at 3 and 2 mm), velocity resolution ($V_{\text{res}}$) of the spectra, main beam efficiency ($B{{}_{\text{eff}}}$), forward efficiency ($F_{\text{eff}}$), and system temperature ($T_{\text{sys}}$).
@@ 3.1 Line fitting
The spectra were fit using the MAdrid Data CUBe Analysis (MADCUBA [^2] MADCUBA is a software developed at the Madrid Center of Astrobiology (CAB, CSIC-INTA) that enables the visualization and analysis of single spectra and data cubes: https://cab.inta-csic.es/madcuba/; martin2019spectral) software. The transitions were identified using the Spectral Line Identification and LTE Modelling (SLIM) tool of MADCUBA, which makes use of the Jet Propulsion Laboratory (JPL; pickett1998submillimeter) and Cologne Database for Molecular Spectroscopy (CDMS; muller2001cologne) catalogs. The lines were fit using the AUTOFIT function of SLIM. This function produces the Gaussian synthetic spectrum that best matches the observed spectrum, assuming local thermodynamic equilibrium (LTE) conditions and the same fitting input parameters for all transitions of a given molecular species, which are total column density, $N_{\text{tot}}$; excitation temperature, $T_{\text{ex}}$; radial systemic velocity of the source, $V_{\text{LSR}}$; full width at half-maximum of the line, FWHM; and angular size of the emission, $\theta_{\text{S}}$.
@@
For the source size, we assumed that the molecular emission fills the telescope beam, except for methanol and formaldehyde, where an estimation of the emission size was made in fontani2022chemoutII and updated in this paper as discussed in Sect. 4.1.1. Such an assumption is justified by the spatial extent of the emission of the dust (see Figs. 4, 5, 6), as well as by the updated H2CO and CH3OH emission sizes shown in Table 8. The excitation temperature, whenever an estimate was not possible through the fit, has been fixed to the one shown in Table 5, which was estimated from the methanol lines for the 15 targets observed in fontani2022chemoutII. For the remaining sources, we adopted an average $T_{\text{ex}}$ of 10.7 K, derived from those sources for which methanol transitions were available. This approximation was necessary because it was not possible to estimate the temperature individually for each source based on the observed transitions. To evaluate the impact of this assumption, we computed $N_{\text{tot}}$ in the range of excitation temperatures derived from methanol ($7\leq T_{\text{ex}}\leq 15\penalty 10000\ K$; see Table 5). In this $T_{\text{ex}}$ range, the densities vary from the values shown in Tables 7 and 8 by a factor of two at most for the species studied (lowest for HCN, HCO+, and H13CO+; highest for SO).
@@
As a detection criterion, we considered a line as detected if the signal-to-noise ratio (S/N) is greater than three (an estimation of the rms noise is given in fontani2022chemout), while the detection is considered as tentative for $2.5\leq\text{S/N}\leq 3$. If a line is undetected (e.g., S/N $\leq 2.5$), MADCUBA provides a $3\sigma$ upper limit estimate to the line’s integrated intensity from Eq. (29) of martin2019spectral. The upper limit on the integrated intensity is converted into a column-density upper limit using the SLIM parameters $T_{\text{ex}}$ and $\theta_{\text{S}}$. For these estimates, the average FWHM for each species across the sample was adopted (Sect. 5.2). This approach assumes that different molecules trace gases with different physical, kinematic, and excitation conditions.
@@
The total column densities derived from transitions observed at frequencies corresponding to beam sizes different from 28′′ (i.e., the largest beam size among the IRAM 30 m telescope spectral windows) have been rescaled by a factor of $(\frac{\theta_{\text{beam}}}{28^{\prime\prime}})^{2}$. For H2CO and CH3OH, the column densities were estimated assuming a specific angular emission size (see Table 8, and Sect. 4.1.1). In cases where the emission size exceeds 28′′, no correction was applied, under the assumption of uniform emission within the beam. Conversely, when the source size is smaller than 28′′, the same scaling factor was applied to account for beam dilution effects. The estimation of the molecular fractional abundances has been performed using
@@ 3.2 Continuum analysis
For this paper, the analysis of these maps is limited to the extraction of the flux density from an angular region of $28\penalty 10000\ ^{\prime\prime}$ centered on the pointing centers of the molecular spectroscopic data. The dimension of the extracted region has been decided to match the larger beam size of IRAM 30 m in the studied spectral windows (Table 1). A complete description and a deeper analysis of the NIKA2 observations will be given in a forthcoming paper (Fontani et al. in prep.).
The column densities of H2, $N_{\text{H}_{2}}$ were derived from the dust-continuum-emission maps. In the scales (and densities) traced by our observations, the thermal dust emission is expected to be optically thin. Under this assumption, we can use the following equation to derive $N_{\text{H}_{2}}$ from the continuum flux density (e.g., battersby2011characterizing):